OpenStudy (lgbasallote):

Note: This is NOT a tutorial. This is a question. \[\LARGE \int _{-1}^{0} \frac{e^{\frac{1}{x}}}{x^3} dx\] please show me the steps :D note: will no longer be making tutorials for an indefinite time due to some legal issues. i will return to doing them once the issues are fixed. thank you

OpenStudy (anonymous):

APPLY ILATE rule

OpenStudy (lgbasallote):

please start from BEGINNING :p

OpenStudy (anonymous):

take \[ 1/x = t\] \[\frac{-dx}{x^2}=dt\] Write x =1/t \[{dx}=\frac{-dt}{t^2}\] Changing the limit, we get Lower limit t =1/-1 = -1 Upper limit t = 1/0 = infinity Substituting this, we get \[\large \int\limits_{-1}^{\infty} \frac{e^t}{\frac{1}{t^3}}{(\frac{-dt}{t^2})}\] \[\large \int\limits_{-1}^{\infty}- e^t tdt\] Now apply udv rule Take u = t dv = e^t dt find du, v from above Now according to udv rule \[\int\limits_{}^{}u dv = u \int\limits_{}^{} dv - v \int\limits_{}^{}du \] Substitute and get your answer :)

OpenStudy (anonymous):

@lgbasallote , is it clear ?? :D

OpenStudy (lgbasallote):

uhmm ill be afk for a while..ill check it later and ask questions :)

OpenStudy (lgbasallote):

ugh i dont get it from beginninng sorry @shivam_bhalla i think it's me :/ i'll just keep practicing